[Brahms-l] Shift report 20050217 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu Feb 17 2005 - 07:54:32 EST
Shifter leader was SB
Shifters were 


1. Beam from previous shift was lost at the beginning of the night shift. RHIC had power supply problem until about 4:15 am. New store since then and running physics. Beam to be dumped at 9:20 am.
2. Pedestal runs: 13213, 13214;
   Setting FS@4 deg. @A3450 MRS@40 deg. @A2255
   RunNo    Trig2     Trig3
   13215  -  251k   -  103k 
   13216  -  202k   -  158k 
   13217  -  221k    - 207k 

3. Checking high voltages, I found that F1-xu and F1-yv on T3 had
tripped may be due to the sudden beam loss at begining of
shift. Trip was cleared by turning them off and back on again.
All cathods and Field1 on T5 had also tripped. Trip was cleared the same way and all is well. 
    


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Feb 17 07:54:48 2005

This archive was generated by hypermail 2.1.8 : Thu Feb 17 2005 - 07:54:56 EST